### v1

# Target Speaker ASR with Whisper

Abstract
We propose a novel approach to enable the use of large, single speaker ASR models, such as Whisper, for target speaker ASR. The key insight of this method is that it is much easier to model relative differences among speakers by learning to condition on frame-level diarization outputs, than to learn the space of all speaker embeddings. We find that adding even a single bias term per diarization output type before the first transformer block can transform single speaker ASR models, into target speaker ASR models. Our target-speaker ASR model can be used for speaker attributed ASR by producing, in sequence, a transcript for each hypothesized speaker in a diarization output. This simplified model for speaker attributed ASR using only a single microphone outperforms cascades of speech separation and diarization by 11% absolute ORC-WER on the NOTSOFAR-1 dataset.
Index Terms:
target-speaker ASR, diarization conditioning, multi-speaker ASR, Whisper
I Introduction
Self-supervised models [1, 2, 3], LLMs [4, 5], and Whisper-style supervised models [6, 7] have demonstrated that scaling up models to use more parameters and extremely large amounts of data can enable the development of accurate automatic speech recognition (ASR) systems, even in relatively challenging environments. However, these models have primarily been used in single-speaker, single-channel ASR systems, whereas most conversations involve multiple talkers and are often recorded by one or more microphones.
Approaches to handle this scenario generally combine multiple systems that perform source separation, speaker segmentation, overlapped speech detection, post-hoc speaker clustering, and ASR in order to produce speaker-attributed conversation transcripts. Alternatively, there are end-to-end systems that transcribe multi-talker speech directly using special tokens or multiple heads [8, 9, 10, 11]. One type of semi-end-to-end system, dubbed target-speaker ASR (TS-ASR) [12, 13, 14, 15, 16], uses the original input mixture and transcribes each speaker separately. Internally, these models rely on source separation to isolate the target speaker’s speech [17, 18].
Another approach to TS-ASR works by learning to transcribe a single target-speaker’s transcript from a mixture of sources by conditioning on a pre-extracted speaker embedding. Typically, this requires a pretrained speaker embedding extractor [19] and training the ASR model from scratch. More recent methods include the use of adaptation layers and soft prompts to modify existing ASR models to work with speaker embeddings [14]. Because these models are often trained on simulated datasets due to the limited availability of multi-talker ASR datasets and the need for a large number of speaker identities, there is signficant performance degradation when such models are deployed on real multi-speaker data “in the wild” [20, 21, 22].
Figure 1: Proposed Diarization-Conditioned Whisper model. An input audio segment with potentially multiple speakers is augmented with frame-level diarization outputs $\begin{bmatrix}p_{S}^{t}&p_{T}^{t}&p_{N}^{t}&p_{O}^{t}\end{bmatrix}^{T}$ for each of the STNO classes at every frame $t$. Affine transformations, indicated as additions to the left of the Whisper model, are applied to intermediate input representations $\mathbf{z}_{n}^{1:T}$ to generate new embeddings, where $n$ stands for the index of the layer. The final frame-level embedding is a convex combination of these embeddings for each frame.
TABLE I: Comparison of the proposed system build with Whisper-large-v3, trained and tested separately on each dataset, alongside various multi-talker ASR systems. The top section includes systems where no additional information about speaker identity or segmentation is provided. The bottom section features models that directly or indirectly utilize segmentation information.
AMI-sdm
NOTSOFAR-1
Libri2Mix
test-both
Kanda et al. [23] 25.8 - -
Raj et al. [8] 44.6 60.9 -
Vinnikov et al. [20] 35.5
Input masking 79.1 76.6 56.7
Proposed 48.5 24.5 17.6
Ma et al. [14] - - 26.4
Zhang et al. [15] - - 23.5
In this paper, we propose a semi-end-to-end approach to TS-ASR that uses Whisper in a new way. Unlike previous TS-ASR methods, our system does not rely on speaker embeddings, but instead conditions directly on frame-level diarization outputs. We believe that, compared to the aforementioned embedding-based approaches, only ”relative” differentiation between speakers is needed; the TS-ASR system does not need to adapt to an existing subspace of speaker embeddings. Training our model on labeled examples of both target and non-target speech, may also improve speaker discrimination and improve robustness to diarization errors.
To validate our approach, we fine-tune Whisper models on the NOTSOFAR-1 [20], AMI [21], and Libri2Mix [24] datasets using ground truth speaker segmentation. During inference, we utilize the ground truth speaker segmentation as well. All experiments conducted in this study adhere to the conditions of the NOTSOFAR-1 Challenge [^1] https://www.chimechallenge.org/current/task2/index.
II Diarization-Conditioned Whisper
This section presents the Diarization-Conditioned Whisper, a model built upon the Whisper architecture, designed to perform TS-ASR by conditioning on frame-level diarization outputs. An overview of the proposed model is shown in Fig. 1. We adapt Whisper for TS-ASR by adding Frame-Level Diarization Dependent Transformations (FDDT) modules, described in Section II-C. These modules transform the model’s internal representations in order to differentiate between the target- and non-target speakers in the audio.
II-A Silence, Target, Non-Target, and Overlap Masks
Let $\mathbf{D}\in[0,1]^{S\times T}$, where $S$ is the number of speakers in the recording, and $T$ is the number of frames. The matrix $\mathbf{D}$ represents the diarization output, with $d(s,t)$ denoting the probability that speaker $s$ is active at time frame $t$.
The dependency on the number of speakers in $\mathbf{D}$ can be a limiting factor for easily incorporating this mask into the model. To address this, let $s_{k}$ represent the target speaker. We define a distribution over the following mutually exclusive events for a frame at time $t$.
  • S: The time frame, $t$, is silence
  • T: The target speaker, $s_{k}$, is the only active speaker at time frame, $t$.
  • N: One or more non-target speakers, $s\neq s_{k}$ is active and the target speaker, $s_{k}$, is not active at the time frame, $t$.
  • O: The target speaker $s_{k}$ is active while at least one non-target speaker $s\neq s_{k}$ is also active at time frame, $t$.
We define the following distribution over these events.
$p_{S}^{t}=p\left(t=S\right)$ $=\prod_{s=1}^{S}(1-d(s,t))$ (1)
$p_{T}^{t}=p\left(t=T\right)$ $=d(s_{k},t)\cdot\prod_{\begin{subarray}{c}s=1\\ s\neq s_{k}\end{subarray}}^{S}(1-d(s,t))$ (2)
$p_{N}^{t}=p\left(t=N\right)$ $=\left(1-p_{S}^{t}\right)-d\left(s_{k},t\right)$ (3)
$p_{O}^{t}=p\left(t=O\right)$ $=d(s_{k},t)-p_{T}^{t}$ (4)
This definition allows us to use a fix-sized STNO (Silence, Target, Non-target, Overlap) mask $\mathbf{M}^{t}=\begin{bmatrix}p_{S}^{t}&p_{T}^{t}&p_{N}^{t}&p_{O}^{t}\end{bmatrix}^{T}$. The mask is speaker-dependent so that different STNO masks will result in different transcripts by the system.
TABLE II: Analysis of different constraints applied to the FDDT parameters and methods used to initialize them evaluated with Whisper-medium.en on NOTSOFAR-1 eval-small. The FDDT parameters column specifies which parameters are used to condition the model. $\mathbf{W}_{diag}=diag(\mathbf{w})$, where $diag(\mathbf{w})$ is a diagonal matrix with elements of $\mathbf{w}$ in the diagonal.
Initilization Method
FDDT parameters Random Identity Suppressive
$\mathbf{b}$ 28.4 28.0 28.0
$\mathbf{W}_{diag},\mathbf{b}$ 129.4 27.3 26.7
$\mathbf{W},\mathbf{b}$ 129.0 46.1 44.6
II-B Input Masking
Having the STNO mask, a straightforward way to perform target speaker ASR is to mask the signal by multiplying each frame by the probability that it is target speech or involves overlap with the target speaker.
Let $\mathbf{X}\in\mathbb{R}^{F\times T}$ denote the matrix of input features, where $F$ is the number of feature dimensions (e.g., mel-filter banks). The masked feature matrix $\mathbf{X}_{\text{masked}}$ is computed as:
$\mathbf{X}_{\text{masked}}(f,t)$ $=\mathbf{X}(f,t)\cdot(p_{T}^{t}+p_{O}^{t}).$ (5)
Here, we add $p_{T}^{t}$ and $p_{O}^{t}$ to ensure that both target speech and overlapping speech are preserved in the masked features. Frames where neither the target speech nor overlap is present are effectively masked out (i.e., set to zero).
However, similar to source separation approaches, this method has limitations. It can introduce artifacts because we are creating a modified version of the input signal, and errors in diarization can propagate through the system, potentially affecting the model’s performance.
II-C Frame-Level Diarization Dependent Transformations
To overcome issues of Input Masking, we designed a soft version called Frame-Level Diarization Dependent Transformations (FDDT). This approach modifies the frame-by-frame model inputs based on the diarization outputs.
Let $\mathbf{Z}^{n}\in\mathbb{R}^{d_{\text{model}}\times T}$ represent the frame-by-frame inputs to $n$-th (transformer) layer. We transform these states by applying four affine STNO layer- and class-specific transformations $\mathbf{W}_{S}^{n},\mathbf{W}_{T}^{n},\mathbf{W}_{N}^{n},\mathbf{W}_{O}^{n}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{model}}}$ together with biases $\mathbf{b}_{S}^{n},\mathbf{b}_{T}^{n},\mathbf{b}_{N}^{n},\mathbf{b}_{O}^{n}\in\mathbb{R}^{d_{\text{model}}}$ to obtain new speaker-specific states $\hat{\mathbf{Z}}^{n}$ as follows:
$\hat{\mathbf{Z}}^{n}$ $=\left(\mathbf{W}_{S}^{n}\mathbf{Z}^{n}+\mathbf{b}_{S}^{n}\right)\mathbf{p}_{S}+\left(\mathbf{W}_{T}^{n}\mathbf{Z}^{n}+\mathbf{b}_{T}^{n}\right)\mathbf{p}_{T}+$
$+\left(\mathbf{W}_{N}^{n}\mathbf{Z}^{n}+\mathbf{b}_{N}^{n}\right)\mathbf{p}_{N}+\left(\mathbf{W}_{O}^{n}\mathbf{Z}^{n}+\mathbf{b}_{O}^{n}\right)\mathbf{p}_{O}.$ (6)
These transformations create four distinct representations of the frame-by-frame inputs, each emphasizing one of the STNO classes. A new, target-speaker specific representation is formed from the convex combination of these 4 terms, where the term weights come from the STNO mask. Note that the same transformation will be applied to all frames with identical STNO masks.
TABLE III: Effect of enlarging the number of encoder layers, where FDDTs are applied on NOTSOFAR-1 eval-small with Whisper-medium.en (24 encoder layers).
Initilization Method
FDDT parameters # layers Random Identity Suppressive
$\mathbf{b}$ 1 28.7 30.9 29.3
12 28.7 27.6 27.6
24 28.4 28.0 28.0
$\mathbf{W}_{diag},\mathbf{b}$ 1 117.7 27.8 27.0
12 118.9 27.4 27.1
24 129.4 27.3 26.7
Including biases in the affine transformations is crucial as it enables the model to differentiate between different STNO types of speech. The biases can shift the representations, making it easier for the model to recognize and distinguish between silence, target speaker speech, non-target speaker speech, and overlapping speech.
On the other hand, by using matrices, $\mathbf{W}_{S,T,N,O}$, we can transform hidden states to the space where it is possible to distinguish between speakers more efficiently or even suppress some parts of the signal.
Fine-tuning the model using randomly initialized FDDT matrices could easily disrupt the internal representations of the model. Therefore, we propose initialization strategies to mitigate this risk:
  • Identity Initialization (Non-Disturbing Init): Here, biases are initialized with zero vectors, and weights are initialized as identity matrices. This method ensures that the model’s internal representations are not altered.
  • Suppressive Initialization: To bias the model toward masking other speakers, we initialize the $\mathbf{W}_{S,N}$ weights as diagonal matrices with values close to zero, e.g., 0.1. This approach helps the model to distinguish between different types of speech, reinforcing the separation between the STNO classes.
TABLE IV: The performance of FDDT given a reduction of information provided from diarization output on NOTSOFAR-1 eval-small with Whisper-medium.en. When employing the STNO mask, all frames are transformed, while with TNO, frames corresponding to the silence are left unchanged.
STNO TNO TN T
26.7 30.0 28.7 34.8
III Experiments
We primarily conducted our experiments on the new NOTSOFAR-1 dataset [20], which includes approximately 315 meetings, each averaging 6 minutes, capturing a broad range of real-world acoustic conditions and conversational dynamics. To verify generalization and demonstrate the competitiveness of the proposed method, we evaluated our best models performance on commonly used synthetic dataset Libri2Mix [24] as well as on real-world meeting recordings like AMI [21] and NOTSOFAR-1. All experiments were conducted in compliance with CHiME8-NOTSOFAR1 rules.
They are divided into two parts. In Section III-C, we examine the behaviour of FDDT under different weight structure constraints, initialization methods, the number of additional parameters, and the information provided. In Section III-D, we analyze the framework’s performance when scaled.
Source codes and recipes [^2] https://github.com/BUTSpeechFIT/TS-ASR-Whisper are built on top of the transformers library [25]. All models are evaluated with the Optimal Reference Combination WER (ORC-WER) [26]. For brevity, we will refer to this metric as WER throughout the rest of the text.
III-A Training details
Adapting the acoustic part of the model without negatively impacting the generalization of the decoder can be challenging, especially with models like Whisper. To address this, we incorporated an additional CTC (Connectionist Temporal Classification) head, following the hybrid CTC-attention-based training scheme proposed in [27]. Given Whisper’s large 50k vocabulary size and fixed sequence length, adding an extra projection layer poses memory challenges. Therefore, we introduced two convolutional layers with a subsampling factor of two each and an additional self-attention layer to optionally realign the sequence. The CTC head can also be used during joint decoding to help mitigate hallucinations [28].
Both the CTC head and the decoder are trained with timestamp tokens. The logits for the CTC blank token are produced by a trainable projection $\mathbf{W}_{\text{blank}}\in\mathbb{R}^{d{{}_{\text{model}}}\times 1}$, and we use $\alpha=0.3$ as the CTC loss weight.
All models are trained with an overall batch size of 64 samples using bf16 precision and the AdamW optimizer [29]. The learning rate is set to $2\times 10^{-6}$, with a weight decay of $1\times 10^{-6}$, a linear decay scheduler, and 2k warm-up steps. The new parameters introduced by FDDT are trained with a learning rate of $2\times 10^{-4}$. By default, FDDT modules are inserted before all layers of the encoder with the diagonal constraint. Unless otherwise stated, the CTC head undergoes an initial ”CTC preheating” phase, where it is trained on LibriSpeech for 10k steps, with the rest of the model frozen. Afterwards, FDDT and CTC parameters are trained for a single epoch (Amplification phase). Finally, the full model is trained for up to 50k steps, with early stopping set to patience of 5 epochs. Most of the models typically converge within ten epochs. For the final evaluation, we always select the best-performing checkpoint based on the development set WER.
TABLE V: Different sizes of training corpuses affecting the performance of Whisper-medium.en. Tested on NOTSOFAR-1 Eval Small.
NOTSOFAR-1 + AMI + Libri2Mix
26.7 25.6 24.8
III-B Baseline Comparison
Table I compares the proposed method with different end-to-end and modular systems. It can be seen that our approach vastly outperforms the naive approach input masking on all three datasets mainly due to finetuning and the ability to handle overlapped speech. Even though our approach does not outperform any baselines in the first part of the table, it out performs the NOTSOFAR baseline [20] and the finetuned SURT model on the NOTSOFAR-1 dataset as well. Lastly, our approach outperforms all other approaches on the synthetic dataset Libri2Mix.
III-C Frame-Level Diarization Dependent Transformation
To evaluate the impact of the Frame-by-Frame Diagonal Transformations (FDDT), we conducted a study to determine whether limiting the number of parameters in the additional modules affects system performance and how important correct initialization is. Table II shows that using biases alone performs similarly to diagonal matrices, indicating that biasing frame-by-frame representations can effectively focus the model on frames corresponding to the same class. The table also highlights that randomly initializing FDDT parameters is suboptimal and can significantly harm model performance, suggesting that suppressive initialization is preferable. Interestingly, using full weights leads to noticeable performance degradation, likely because it disrupts the frame-by-frame representations. Table III demonstrates that even a single layer of bias-only parameters can achieve performance comparable to the best diagonal setup.
Table IV further shows that a model can perform TS-ASR with just a single STNO class corresponding to the target frames without significant performance degradation. Interestingly, when using three classes (TNO), the model performs worse than when using STNO or TN, which is counterintuitive since the model theoretically receives the same amount of information.
III-D Scaling System With More Data And Parameters
Given that even a simple approach using a single bias performs well, Table V demonstrates the improvement in system performance when additional training data is used. The results highlight that incorporating synthetic data provides additional gains beyond those achieved with real meeting data alone, raising the question of whether pretraining on synthetic data might offer further improvements.
Table VI provides a performance analysis across different model sizes, highlighting the improvements gained from employing an additional CTC head. The analysis also explores the impact of CTC preheating and an Amplification phase, showing how these techniques can optimize performance across varying model scales.
TABLE VI: Influence of CTC head and size of the model evaluated on NOTSOFAR-1 eval-small.
small.en medium.en large-v3
without CTC 30.3 28.1 24.6
with CTC 31.0 28.9 25.8
+ CTC Preheating 29.2 26.7 25.2
+ Amplification phase 30.3 27.4 24.5
IV Conclusions and Limitations
In this study, we have demonstrated the efficacy of our approach and analyzed its setup across multi-domain datasets. However, further validation with additional datasets, especially in different conditions or languages, is necessary to confirm its generalizability and scalability.
Our analysis has highlighted both strengths and limitations of the model. Nevertheless, a more in-depth examination of errors and performance under specific conditions is essential to understand and fully optimize the system. Future work will analyze these errors and explore strategies to improve the model’s robustness against diarization errors.
The investigation into the usage of synthetic data suggests that training a TS-ASR model from scratch using only synthetic data with learnable embeddings, followed by fine-tuning on target data, is also a promising avenue. It remains to be seen whether this approach will achieve the same level of performance as models like Whisper, which benefit from extensive pre-training on diverse datasets.
Moreover, our method can be applied to other pre-trained ASR models. Integrating diarization information into these models using our approach could provide valuable insights into its versatility and effectiveness. Comparing the performance of TS-ASR across various ASR architectures will be an important step in evaluating its adaptability and benefits.